OpenStudy (anonymous):

Lecture #9 homework check. Yet again, most of the textbook questions do not have answers in the back of the book. Yet again, I want to check my answers to see if I royally messed up.

OpenStudy (anonymous):

A&E 7-19)\[\text{a) False, }IE \propto \chi\]\[\text{b) True, electrons are less tightlly held.}\]\[\text{c) True, }IE \propto \chi, \space EA \propto \chi\]\[\text{d) True, the same general trend is seen for } \chi\]\[\text{e) True, electronic configuration determines }\chi\]\[\text{f) True, as a general trend, IE increases from right to left across a row}\] A&E 7-28) Not all elements form stable covalent bonds. A&E 8-88) The Lewis dot structure assumes atoms strive for octet stability. Hypervalency is not really accounted for. \[\text{A&E 8-106) NO} \rightarrow covalent; \space \overrightarrow{HF} \rightarrow polar\space covalent; \space \overrightarrow{MgO}\rightarrow ionic;\]\[\overrightarrow{AlCl_3}\rightarrow polar \space covalent; \space \overrightarrow{SiO_2} \rightarrow polar \space covalent; \space \overrightarrow{C=O} \rightarrow polar \space covalent\] \[\text{A&E 8-110) }Q=\frac{\mu}{r}=\frac{1.826D}{92pm}=6.621(10^{-20})C\approx\text{41.3% ionic character}\] \[\text{A&E 8-112) }Q=\frac{\mu}{r}=\frac{3.59D}{228.69pm}\approx\text{32.68% ionic character for PbS(g)}\]\[Q=\frac{\mu}{r}=\frac{4.64D}{192.18pm}\approx\text{50.27% ionic character for PbO(g)}\]PbS(s) is covalent; PbO(s) is polar covalent. \[\text{A&E 9-71) } \overrightarrow{NO}\text{(very slight net); }\overrightarrow{HF};\space \overrightarrow{PCl_3};\space\overrightarrow{CO_2}\text{(none net); }\overrightarrow{SO_2}\]